What is the pH of aqueous 0.10 M HNO2? K (HNO3) = 4.5 * 10-4 O a. 1.67 Ο 6. 3.35 ι Ο υ τ Ο ο Ο υ
Answer #1

HNO2 = H+ + NO2-

I 0.1 0 0

C -X   +X +X

E (0.1-X) +X +X

Ka = 4.5 x 10-4

Ka = (H+)(NO2-)/(HNO2)

4.5 x 10-4 = X2 / (0.1-X)

X = -0.0069 or 0.0069

So H+ = 0.0069

pH = -log(H+)

pH = -log (0.0069)

pH = 2.1611 = 2.17

so option C is correct